OpenStudy (anonymous):

Factor the GCF first and then factor the trinomial: 2x2 + 20x + 50 A. 2(x + 5)(x + 5) B. 2(x + 5)(x + 10) C. 2(x + 2)(x + 25) D. prime Factor the GCF first and then factor the trinomial: 2x2 + 20x + 50 A. 2(x + 5)(x + 5) B. 2(x + 5)(x + 10) C. 2(x + 2)(x + 25) D. prime @Mathematics

OpenStudy (saifoo.khan):

D

OpenStudy (anonymous):

A. 2(x + 5)(x + 5)

OpenStudy (anonymous):

A

OpenStudy (anonymous1):

\[2x^2+20x+50=2(x^2+10x+25)\] Now factor \[x^2+10x+25\] if it is a perfect square then it has the form \[(a+b)^2=(a+b)(a+b)==a^2 + 2ab+b^2\] Taking a=x and b=5: \[(x+5)^2=(x+5)(x+5)=x^2+2*5*x+5^2\] Thus: \[2x^2+20x+50=2(x^2+10x+25)=2(x+5)^2=2(x+5)(x+5)\] Letter A. 2(x + 5)(x + 5).